OpenStudy (anonymous):

The cost of 6 movie tickets is $57.00. What is the cost of 8 movie tickets? A. $59.00 B. $63.00 C. $65.00 D. $76.00

OpenStudy (anonymous):

@CGGURUMANJUNATH @miyah04 @madyhughes @Mindblast3r @Squirrels @Qordix @quirah @DaniSchwarz

jimthompson5910 (jim_thompson5910):

"The cost of 6 movie tickets is $57.00" so what is the cost of one movie ticket?

OpenStudy (anonymous):

do i have to divide ?

jimthompson5910 (jim_thompson5910):

yes, what do you get?

OpenStudy (anonymous):

9.5

OpenStudy (anonymous):

its d if u divide 57 by the 6 tickets equals 9.5 then multiply by 8 you get 76

jimthompson5910 (jim_thompson5910):

so the cost of 1 ticket is $9.50 now multiply that by 8 to get the cost of 8 tickets

OpenStudy (anonymous):

thanks it was confusing but i get it now

OpenStudy (texaschic101):

or this way...a proportion 6/57 = 8/x --- (6 tickets to $57 = 8 tickets to x) cross multiply (6)(x) = (57)(8) 6x = 456 x = 456/6 x = 76 8 tickets cost $76

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

do it whichever way is easiest for you :)